# Real Time Semantic Segmentation of High Resolution Automotive LiDAR Scans

### v1

@@ III-A Spherical Projection
Our approach relies on a staggered image-like representation, where neighboring pixels correspond to adjacent LiDAR measurement rays and their respective measured points, as introduced in [20]. This representation is obtained through a spherical image projection. Each point in the point cloud, defined by its Cartesian coordinates $[x,y,z]^{T}$, is transformed into spherical coordinates, represented as $[\phi,\theta,r]^{T}$. The azimuth $\phi$ corresponds to the angle of the point in the $xy$-plane, the inclination $\theta$ denotes the angle from the positive $z$-axis, and $r$ is the distance from the origin. This spherical projection captures the geometry of the sensor in a single image. We then use the following projection model to obtain a pixel $\vec{u}$ of a 3D point in a staggered spherical image representation:
@@ IV Experiments
CENet [9] HarDNet 6.8M 63ms 90.29 55.86 71.66 58.54 92.44 85.86 57.45 41.15 39.53 52.51
FIDNet [8] ResNet34 6.05M 66ms 91.86 66.18 69.00 59.24 91.17 84.06 41.40 49.64 57.57 53.58
Our ResNet18 14.17M 8ms 86.45 62.98 74.62 59.41 91.72 83.38 39.14 43.04 52.25 53.57

### v2

@@ Abstract
©2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ III-A Spherical Projection
Our approach relies on a staggered image-like representation, where neighboring pixels correspond to adjacent LiDAR measurement rays and their respective measured points, as introduced in [20]. This representation is obtained through a spherical image projection. Each point in the point cloud, defined by its Cartesian coordinates $x,y,z]^{T}$, is transformed into spherical coordinates, represented as $\phi,\theta,r]^{T}$. The azimuth $\phi$ corresponds to the angle of the point in the $xy$-plane, the inclination $\theta$ denotes the angle from the positive $z$-axis, and $r$ is the distance from the origin. This spherical projection captures the geometry of the sensor in a single image. We then use the following projection model to obtain a pixel $\vec{u}$ of a 3D point in a staggered spherical image representation:
@@ IV Experiments
CENet [9]
HarDNet 6.8M 63ms 90.29 55.86 71.66 58.54 92.44 85.86 57.45 41.15 39.53 52.51
FIDNet [8]
ResNet34 6.05M 66ms 91.86 66.18 69.00 59.24 91.17 84.06 41.40 49.64 57.57 53.58
Our
ResNet18 14.17M 8ms 86.45 62.98 74.62 59.41 91.72 83.38 39.14 43.04 52.25 53.57